# A Prompt Learning Framework for Source Code Summarization

### v1

@@ 5.2.1. RQ1: Effectiveness of PromptCS
5.2.6. RQ6. PromptCS’s performance in human evaluation
In addition to automated evaluation, we conduct a human evaluation by following the previous works (Wu et al., 2021; Hu et al., 2020; Wei et al., 2020; Zhang et al., 2020; Hu et al., 2022a) to evaluate the summaries generated by three baselines and PromptCS. Specifically, we invite five volunteers with more than four years of software development experience and excellent English ability to carry out the evaluation. We randomly select 100 code snippets from the CSN-Java dataset, the corresponding ground-truth summaries, and summaries generated by baselines and PromptCS. Each volunteer is asked to assign scores from 1 to 5 to the generated summaries based on how similar they are to the corresponding ground-truth summaries, where 1 means “Not Similar At All” and 5 means “Highly Similar/Identical”. To ensure the fairness and objectivity of experimental results, each summary is evaluated by five volunteers, and the final score is the median value of their scores.
Table 8. Results of human evaluation.
@@
Table 8 shows the score distribution of the generated summaries. Observe that PromptCS achieves the best scores and improves the average (Avg.) score from 2.31 (instruction prompting with zero-shot learning), 3.31 (instruction prompting with few-shot learning), and 3.60 (task-oriented fine-tuning) to 3.64. Specifically, among the randomly selected 100 code snippets, PromptCS can generate 15 highly similar or even identical summaries with the ground-truth ones (score = 5), 56 good summaries (score $\geq$ 4) and 93 summaries that are not bad (score $\geq$ 3). PromptCS also receives the smaller number of negative results (score $\leq$ 2). Based on the 100 final scores for each baseline and PromptCS, we follow (Zhang et al., 2020) and conduct Wilcoxon signed-rank tests (Wilcoxon et al., 1963) and compute Cliff’s delta effect sizes (Macbeth et al., 2011). Comparing PromptCS with instruction prompting with zero-shot learning, instruction prompting with few-shot learning, and task-oriented fine-tuning, the p-values of Wilcoxon signed-rank tests at 95% confidence level are 5.46E-16, 0.016, and 0.722, which means the improvements achieved by PromptCS are statistically significant over instruction prompting with zero-shot and few-shot learning. In addition, Cliff’s delta effect sizes are 0.5669 (large), 0.1478 (small), and 0.0314 (negligible), respectively.
@@ 6. Discussion
Figure 6. Code summarization case 1
@@
Figure 7. Code summarization case 2
@@
7. Related Work
7.1. Code Summarization
@@
Research on code summarization started more than a decade ago. Most early code summarization techniques (Haiduc et al., 2010a; Sridhara et al., 2010; Haiduc et al., 2010b; Moreno et al., 2013) are extractive methods. Such methods work by extracting a subset of the statements and keywords from the code, and then including information from those statements and keywords in the final generated summary. For instance, Sonia Haiduc et al. (Haiduc et al., 2010a) first propose an extractive method to generate summaries for code snippets. Subsequently, in light of the widespread availability of extensive code datasets and the success of neural machine translation (NMT) research in the field of NLP (Cho et al., 2014a; Bahdanau et al., 2015), deep learning approaches have increasingly become prevalent in enhancing the capabilities of code summarization tools. Pioneers in the field have discovered that certain Seq2Seq models, including RNN (Cho et al., 2014b) and LSTM (Iyer et al., 2016), demonstrate the ability to effectively capture the semantic relationships between code snippets and their corresponding summaries. For example, Iyer et al. (Iyer et al., 2016) employ LSTM networks with attention to generating sentences describing C# code snippets and SQL queries. Hu et al. (Hu et al., 2018a) combine SBT with LSTM to automatically generate code comments for Java methods. However, it is noteworthy that RNN-based methodologies encounter challenges in maintaining long-term dependencies. To overcome this limitation and address potential bottlenecks in modeling long sequences, some Transformer-based methods have been proposed. Zhang et al. (Zhang et al., 2020) propose a retrieval-based neural source code summarization approach by enhancing a Transformer-based model with the most similar code snippets retrieved from the training set. Ahmad et al. (Wu et al., 2021) explore the effectiveness of using Transformer to capture long-range dependencies in code for generating code summaries. In this paper, we focus on adapting LLMs for code summarization tasks (Ahmed and Devanbu, 2022b; Sun et al., 2023c; Tian et al., 2023).
7.2. LLM for SE
Over the years, developers have frequently sought programming assistance during the development process through channels such as blogs and documentation. This involves acquiring code and adapting it to their own context, as well as accessing explanations for related code to facilitate comprehension.
Nowadays, developers can also leverage LLMs for such tasks. GPT-Neo (Black et al., 2021), produced by EleutherAI in 2021, is an autoregressive LLM primarily designed for generating natural language text but can also be applied to generate source code. PolyCoder (Xu et al., 2022), released in 2021, features 2.7B parameters and is trained on 249GB of code across 12 programming languages. In 2022, Salesforce introduced a language model called CodeGen (Nijkamp et al., 2023), specifically designed for code generation tasks. It employs a multi-step program synthesis approach, where a single program is factorized into multiple prompts specifying subproblems. That same year, Facebook proposed the first generative model InCoder (Fried et al., 2023), capable of directly performing zero-shot code infilling. In 2023, the BigCode community presented StarCoder (Li et al., 2023), a 15.5B model supporting over 80 programming languages. It boasts infilling capabilities and fast large-batch inference enabled by multi-query attention.
8. Conclusion
@@
Acknowledgment
The authors would like to thank the anonymous reviewers for their insightful comments. This work is supported partially by National Natural Science Foundation of China (61932012, 62141215, 62372228).

### v2

@@ 4.2.1 RQ1: Effectiveness of PromptCS
4.2.6 RQ6. PromptCS’s performance in human evaluation
In addition to automated evaluation, we conduct a human evaluation by following the previous works [70, 46, 76, 11, 77] to evaluate the summaries generated by three baselines and PromptCS. Specifically, we invite five volunteers with more than four years of software development experience and excellent English ability to carry out the evaluation. We randomly select 100 code snippets from the CSN-Java dataset, the corresponding ground-truth summaries, and summaries generated by baselines and PromptCS. Each volunteer is asked to assign scores from 1 to 5 to the generated summaries based on how similar they are to the corresponding ground-truth summaries, where 1 means “Not Similar At All” and 5 means “Highly Similar/Identical”. To ensure the fairness and objectivity of experimental results, the order of the summaries in each case is shuffled and each summary is evaluated by five volunteers, and the final score is the median value of their scores.
Table 8: Results of human evaluation.
@@
Table 8 shows the score distribution of the generated summaries. Observe that PromptCS achieves the best scores and improves the average (Avg.) score from 2.31 (instruction prompting with zero-shot learning), 3.31 (instruction prompting with few-shot learning), and 3.60 (task-oriented fine-tuning) to 3.64. Specifically, among the randomly selected 100 code snippets, PromptCS can generate 15 highly similar or even identical summaries with the ground-truth ones (score = 5), 56 good summaries (score $\geq$ 4) and 93 summaries that are not bad (score $\geq$ 3). PromptCS also receives the smaller number of negative results (score $\leq$ 2). Based on the 100 final scores for each baseline and PromptCS, we follow [11] and conduct Wilcoxon signed-rank tests [78] and compute Cliff’s delta effect sizes [79]. Comparing PromptCS with instruction prompting with zero-shot learning, instruction prompting with few-shot learning, and task-oriented fine-tuning, the p-values of Wilcoxon signed-rank tests at 95% confidence level are 5.46E-16, 0.016, and 0.722, which means the improvements achieved by PromptCS are statistically significant over instruction prompting with zero-shot and few-shot learning. In addition, Cliff’s delta effect sizes are 0.5669 (large), 0.1478 (small), and 0.0314 (negligible), respectively.
@@ 5 Case Study
Figure 7: Code summarization case 1
@@
Figure 8: Code summarization case 2
@@
6 Related Work
6.1 Code Summarization
@@
Research on code summarization started more than a decade ago. Most early code summarization techniques [8, 80, 81, 82] are extractive methods. Such methods work by extracting a subset of the statements and keywords from the code, and then including information from those statements and keywords in the final generated summary. For instance, Sonia Haiduc et al. [8] first propose an extractive method to generate summaries for code snippets. Subsequently, in light of the widespread availability of extensive code datasets and the success of neural machine translation (NMT) research in the field of NLP [47, 48], deep learning approaches have increasingly become prevalent in enhancing the capabilities of code summarization tools. Pioneers in the field have discovered that certain Seq2Seq models, including RNN [83] and LSTM [67], demonstrate the ability to effectively capture the semantic relationships between code snippets and their corresponding summaries. For example, Iyer et al. [67] employ LSTM networks with attention to generating sentences describing C# code snippets and SQL queries. Hu et al. [10] combine SBT with LSTM to automatically generate code comments for Java methods. However, it is noteworthy that RNN-based methodologies encounter challenges in maintaining long-term dependencies. To overcome this limitation and address potential bottlenecks in modeling long sequences, some Transformer-based methods have been proposed. Zhang et al. [11] propose a retrieval-based neural source code summarization approach by enhancing a Transformer-based model with the most similar code snippets retrieved from the training set. Ahmad et al. [70] explore the effectiveness of using Transformer to capture long-range dependencies in code for generating code summaries. Recently, there are also a lot of studies that apply LLMs to code summarization tasks [84]. For example, Ahmed et al. [28] investigate the effectiveness of few-shot training in adapting LLMs to code summarization and find that it can make Codex significantly outperform fine-tuned pre-trained language models, such as CodeT5. Given the concern of potential code asset leakage when using commercial LLMs (e.g., GPT-3.5), Su et al. [85] utilize knowledge distillation technique to distill small models from LLMs. Their experimental findings reveal that the distilled small models can achieve comparable code summarization performance to LLMs. Gao et al. [30] investigate the optimal settings for in-context learning, including few-shot example selection methods, few-shot example order, and the number of few-shot examples. They find that carefully designed few-shot examples can significantly improve LLMs’ performance on code-related tasks (including code summarization tasks). Geng et al. [86] reveal the ability of LLMs to address multi-intent comment generation. Ahmed et al. [87] propose to enhance few-shot samples with semantic facts automatically extracted from the source code. Sun et al. [5] design some heuristic questions to collect the feedback of ChatGPT, thereby finding an appropriate prompt to guide ChatGPT to generate in-distribution code summaries. In our work, we train a prompt agent under the guidance of LLMs to generate continuous prompts. Our method is non-invasive to LLMs and freezes the parameters of LLMs when training the prompt agent, which can greatly reduce the requirements for training resources.
6.2 LLM for SE
With the success of LLMs in NLP, a large number of SE researchers have started integrating them into the resolution process of various SE tasks [88, 19]. For example, Han et al. [89] leverages LLMs combined with in-context learning to interpret software requirements from textual descriptions for code generation. Jain et al. [90] explore the integration of LLMs (e.g., Codex) in generating code from natural language descriptions of programmer intent. Dakhel et al. [91] utilize mutation testing to augment the prompts, guiding LLMs to generate test cases that can detect software bugs. Sun et al. [92] attempt to detect smart contract logic vulnerabilities by combining GPT and static analysis. In this paper, we focus on adapting LLMs for code summarization tasks [28, 5, 15].
Adapting LLMs to SE with PEFT. As mentioned in Section 1, SE researchers have conducted many empirical studies to investigate the effectiveness of PEFT methods from NLP in adapting LLMs to code-related tasks. For instance, Wang et al. [43] conduct an empirical study to evaluate the usage and effect of prompt tuning in code intelligence tasks. They conduct prompt tuning on CodeBERT and CodeT5 and experiment with three code intelligence tasks including defect prediction, code summarization, and code translation. The intuition of prompt tuning is to convert the training objective of downstream tasks into a similar form as the pre-training stage, i.e., the MLM objective [93, 56]. They find that prompt tuning outperforms fine-tuning and is more effective in low-resource scenarios. Martin et al. [42] investigate the effectiveness of PEFT techniques in automated code generation. They find that LLMs with PEFT consistently and significantly outperform small language models under the same GPU limit. Liu et al. [40] examine the effect of two PEFT methods, namely Adapter Tuning (AT) [94] and Low-Rank Adaptation (LoRA) [37] on code-change-related tasks, including just-in-time defect prediction [95] and commit message generation [96]. Liu et al. [41] conduct an empirical study of PEFT methods (including Adapter [97], LoRA [37], Prefix tuning [36], and MHM [98]) in various code understanding and code generation tasks. Zou et al. [99] examine the effectiveness of five PEFT methods, including Houlsby (i.e., Adapter [97]), Pfeiffer [100], Parallel (i.e., MHM [98]), Prefix tuning, and LoRA, on eight models and four code-related tasks, including clone detection, defect detection, code search, and code translation. Zhuo et al. [101] investigate 5 tasks and 8 different datasets encompassing both code comprehension and code generation tasks on StarCoder [24]. They find that PEFT generally leads to the best downstream performance, and PEFT methods differ significantly in their efficacy based on the model scale. Although these empirical studies cover various PEFT methods and a range of code-related tasks, they merely report the effectiveness of applying PEFT methods to adapt LLMs for code-related tasks. They lack in-depth analysis of the performance of specific PEFT methods and the factors influencing their effectiveness on particular code-related tasks. Unlike them, our PromptCS is an effective and customized prompt learning framework for the code summarization task. In addition, we delve deeply into the factors that may impact PromptCS’ performance, including prompt encoder architecture design, prompt length, and concatenation modes.
7 Conclusion
@@
[…]
The authors at Nanjing University were supported, in part by the National Natural Science Foundation of China (61932012 and 62141215). The authors at Nanyang Technological University are supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008).